How many liters of gas can be collected at 25 C and 0.98atm if 0.063 grams of zinc metal fully reacts in a solution of hydrochloric acid. A 15.5% error was calculate for this reaction, what was the actual volume collected in this experiment?
1
Expert's answer
2020-04-14T01:43:11-0400

T=250C

P=0.98atm

if 0.063g react then 0.63/65.4=0.0096moles react to give 0.0096 moles of H2 gas

Zn+2HCl = ZnCl2 + H2

"V=nRT\/P"

V=(0.0096*8.314*298)/0.98= 24.27m3

1000L=1m3

V=24270L